Run Meeting: August 19th, 2020

From clas12-run
Jump to: navigation, search

Back to Run Group F | Daily Run Group F Coordination Meetings

12:00PM RC Meeting (https://bluejeans.com/7576835804)

RC: M. Hattawy

2020

March
19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30,31
July
31
August
1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20 21 22 23 24, 25, 26, 27, 28, 29, 30, 31
September
1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20 21


Summary of Aug. 19th

Owl shift

  1. Shift summary by the shift expert: https://logbooks.jlab.org/entry/3830429
  2. Issues: RTPC tripped at the beginning of the shift. Beam went down for few hours overnight because of magnets in Acc.
  3. Recorded data: 200nA on 67psig D2: 30M triggers.

Day shift

  1. Shift summary by the shift expert: https://logbooks.jlab.org/entry/3830654
  2. Issues: the scalers and the DC occupancy gets higher than before using the same beam current and target.
  3. Recorded data: 20M events on D2, and 7M on empty target.

Swing shift

  1. Shift summary by the shift expert: https://logbooks.jlab.org/entry/3830951
  2. Issues: Mohammad: done target scan using 100 nA beam -> target center on 2H01 is (0, -0.5) mm in (x,y). The scalers got slightly higher after centering the beam on the target. Done harp scans on 2H01 and 2C24 with 25 nA beam current and the beam profile looks extremely satisfying, looks even better than the reference harp scans. Changed the plan to keep taking data using 200 nA as we already hitting the 6% occupancy in DC region 1.
  3. Recorded data: 32M on full D2 target so far ...


Whole day summary

  1. Total recorded data today: 89M events.
  2. Part-II total accumulated data on D2: 687M.
  3. Concerns: We need to solve the puzzle of the increased scalers and DC occupancy with the new trigger file (we may go back to older trigger files).

Recorded data numbers

Number of events recorded